OpenStudy (anonymous):

What is the first step in solving a quadratic equation of the form given below? (ax + b)2 = c A. Divide both sides by c B. Use the zero product rule C. Factor out a common factor D. Take the square root of both sides

OpenStudy (anonymous):

Any idea?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

\[c = (ax+b)^2\] just curious what are we solving for :o

OpenStudy (anonymous):

But anyways, take a stab at this, think for a bit what you would do to isolate x I guess, first step?

OpenStudy (anonymous):

A or C?

OpenStudy (anonymous):

Good guess, so you know for sure we can eliminate B and C, they might work in a parallel universe with different algebra rules haha. Also, solving for quadratic equation here they mean = 0 or solving for x?

OpenStudy (anonymous):

Well it seems reasonable to solve for x, so can we eliminate anything else?

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

What would make the "squareroot" on the left side = 1, using exponents

OpenStudy (anonymous):

Mhm not quite

OpenStudy (anonymous):

i meant eliminate A

OpenStudy (anonymous):

\[c = (ax+b)^2 \implies \sqrt{c} = \sqrt{(ax+b)^2} \implies \sqrt{c} = (ax+b)^{2/2} \implies \sqrt{c} = (ax+b)\] phew I need some sleep

OpenStudy (anonymous):

Yes, so our answer is D

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no it was right lol thank you

OpenStudy (anonymous):

Oh ok lol, ignore this than, the wording confused me!

OpenStudy (anonymous):

Thank you Lewis

OpenStudy (anonymous):

no problem